# C2SaferRust: Transforming C Projects into Safer Rust with NeuroSymbolic Techniques

### v1

@@ 1 Introduction
Another challenge when applying LLMs for code translation is the length of input and/or output code. LLMs operate on a fixed context window, which means that they can only process inputs and generate outputs that are less than a certain predefined length. A real-world code project with dozens of files and hundreds of lines per file may not fit within this window. Further, even if the entire project does fit, it is known that LLMs tend to produce more errors as the length of the input or output increases. For these reasons, most LLM-based code translation research has focused on programs that are relatively short standalone files, such as those used in competitive programming contests.
@@
Summary of Results: We evaluate our tool on two datasets - one contributed by us, and one from a prior work, Laertes [8].
@@
    The ecosystem of static analysis tools for Rust is rather limited, and so we had to implement most of our analyses as Rust compiler passes, directly using the Rust compiler’s intermediate representations. This lays the groundwork to build future Rust static analysis tools. For instance, we build program slicing and callgraph analysis algorithms, which are quintessential static analysis techniques [14]. We make all of this code publicly available on GitHub [^1] https://github.com/vikramnitin9/c2saferrust.
@@ 2.1 Safety and Unsafety in Rust
Rust is a systems language providing a safe alternative to C and C++ (which are susceptible to bugs and security vulnerabilities [7, 15, 23]) with a robust ownership system ensuring memory and thread safety without requiring garbage collection.
In order to adhere to these safety guarantees, every value in the memory is tied to the variable that owns it, and the memory is automatically de-allocated when the owning variable is no longer in scope. While ownership is exclusive to a variable, Rust also permits other variable to temporarily borrow the values. However, each borrow is further associated with a lifetime while the Rust compiler tracks their corresponding constraints. For instance, it is idiomatic for functions to borrow their arguments from their caller (for the lifetime of the function call) before restoring ownership back to the caller after the call terminates.
Rust further introduces and enforces mutable borrows and immutable borrows which are governed by the principle of exclusivity to ensure memory safety and prevent data races. An immutable borrow (&T) allows read-only access to a value and permits multiple immutable borrows to coexist simultaneously, as they do not interfere with each other. However, while immutable borrows are active, no mutable borrow of the same value is allowed. In contrast, a mutable borrow (&mut T) provides exclusive, writable access to a value, meaning no other borrows—neither mutable nor immutable—can exist simultaneously. This exclusivity ensures that data cannot be read or modified concurrently, maintaining consistency and eliminating potential race conditions. At compile time, a borrow checker enforces these rules, guaranteeing safe and predictable behavior in all borrowing scenarios [27].
The strong safety guarantees in Rust come with several challenges, namely: a) difficulty in implementing complex data structures, such as graphs, where cyclic references or multiple ownerships are common, due to strict ownership and borrowing rules; b) performance bottlenecks such as overhead from bounds checks on array access or runtime borrow checks for dynamically mutable structures; and c) complexities in integrating with system libraries or for writing low-level kernel or driver code via foreign function interface (FFI), which lack Rust’s safety guarantees [3]. To overcome these limitations, which may be too restrictive for several scenarios, Rust provides the unsafe keyword to enable developers to perform operations that the compiler may otherwise not be able to guarantee as being safe. Unsafe blocks (and functions) allow actions such as dereferencing raw pointers, accessing mutable global variables, calling unsafe functions (e.g., through FFI or compiler intrinsics), among other options [26]. While these features are crucial for tasks like interfacing with system libraries, optimizing performance-critical code, or managing complex data structures, they require developers to take full responsibility for ensuring memory safety. Developers who use unsafe must carefully design and test their code to prove its safety, as the compiler’s guarantees no longer apply in these contexts.
The ability to strategically employ unsafe blocks and/or unsafe functions accompanied by strong safety guarantees makes Rust a compelling choice for systems programming. Consequently, there has been a concerted effort to migrate legacy codebases written in C and C++ Rust, leveraging its safety features while eliminating sources of common vulnerabilities that arise due to manual memory management and unchecked behavior [23]. However, this migration process is not straightforward. C and C++ both rely on a very permissive memory models with use of raw pointers, unchecked array indexing, and general lack of enforced ownership or borrowing rules, all of which make manual translation a very arduous and error prone process [16].
@@ 2.2 Translating C to Rust using Transpilation
One of the most direct approaches to translate C code to Rust is to make use of source-to-source compilers (also known as Transpilers). Transpilers aim to translate code written in one programming language to another while endeavoring to preserve the semantic equivalence between the two. The resulting code is intended to be functionally identical to the original. In the context of translating from C to Rust, transpilers such as C2Rust [10] use carefully designed rules to map C constructs into their Rust counterparts.
Challenges with transpilers.  Source to source compilers, such as C2Rust [10], provide an automated approach to translate C to functionally equivalent Rust. However, as discussed previously, transpilers rely on a predetermined set of rule to map constructs like raw pointers, global variables, and manual memory management into their corresponding Rust equivalents. However, these mappings often result in code that compiles in Rust but compromises its safety guarantees, requiring the liberal use of unsafe blocks or foreign functions to maintain functional equivalence. This highlights the difficulty in transforming C’s permissive programming model to that of Rust’s strict ownership and borrowing models. For instance, to dereference raw pointers to conduct pointer arithmetic, C’s pointers are converted Rust’s *const or *mut and encapsulated unsafe blocks, thereby reintroducing potential memory errors. Similarly, global variables in C become static mut in Rust, accessed through unsafe blocks, risking race conditions or undefined behavior in multi-threaded contexts. Unsafe operations like transmute further bypass Rust’s type system, leading to potential alignment errors or undefined behavior. Additionally, C2Rust-transpiled code often retains low-level constructs like malloc and free instead of leveraging idiomatic Rust abstractions like Rc, Arc, or Box, increasing reliance on unsafe and requiring developers to manually ensure memory safety.
@@ 2.3 Programming language translation with Large Language Models (LLMs)
LLMs offer a new paradigm for programming language translation due to their ability to understand and generate code with contextual awareness while reflecting on their training corpus to adhere to recommended safety coding standards [19, 29, 13]. These features enable LLMs to effectively translate programs between languages. Preliminary studies show that while LLMs perform well on crafted benchmarks such as CodeNet [20], Avatar [1], and EvalPlus [17], they struggle to translate complete software repositories, often failing to capture the complexity and inter-dependencies within real-world projects [13]. This limitation arises from challenges such as ensuring syntactic correctness, maintaining semantic equivalence, and adapting to language-specific features. These shortcomings can lead to issues like incorrect API mappings, type mismatches, and functional errors, underscoring the need for enhanced techniques to improve the reliability and accuracy of LLM-based translations [19].
@@ 3 Methodology
In this section, we describe our methodology for translating C code to their safer Rust counterpart, focusing on preserving the intent of the original codebase, but making it more safe. Our tool is designed to operate on complex, real-world C projects, often comprising thousands of lines of code and tens of interdependent files. The translation process prioritizes functional correctness, ensuring that the resulting Rust code behaves as expected under the test cases provided by the user. This section describes our methodology in detail.
@@ 3.1 Overview
C2SaferRust is a neuro-symbolic framework designed to translate C code into safer Rust. We leverage on-demand static analysis and LLMs for translation, and differential testing for validating the translated code. However, given the size of typical C project, directly translating an entire C project to Rust in one step is infeasible. We need to come-up with a strategy to translate and validate smaller pieces of code in a step-by-step manner. Thus, we need a framework that allows partial C and partial (translated) Rust code to coexist, as we need to validate the accuracy of the translation by running this hybrid code. To this end, we adopted the C2Rust framework, which transpiles a given C code repository into an equivalent unsafe Rust repository. While the C2Rust-transpiled code retains most of C’s ‘unsafe‘ properties and original intent, it effectively acts as a Rust wrapper, enabling the unsafe code to execute within a Rust environment. Our translation process begins from this starting point, progressively transforming the unsafe Rust into safer Rust.
This unsafe Rust code is then divided into smaller units, called translation units, using program graphs extracted through symbolic reasoning. Each unit is translated into safer Rust using an LLM, which generates safer code while symbolic reasoning ensures accuracy. Compilation and test cases identify errors, prompting targeted analysis to refine context and guide effective repairs. This iterative approach balances efficiency and accuracy, progressively enhancing translation quality to suit the structure and requirements of each project. An overview of the system is shown in Figure 1. C2SaferRust consists of the following components:
@@
    Translation Orchestrator (§ 3.2): This component, aided by static analysis, orchestrates translation. It breaks the project into smaller translation units that fit within the LLM’s context. It then selects translation units in the optimal order for translation. Translation units are not independent - when one unit is translated, the side effects have to be propagated to all other affected units. The optimal order for translation is moving up the chain of propagated side effects, i.e., translating a unit before translating the units to which its side effects are propagated. This ensures that while the LLM is translating a unit, it can also assimilate the side effects propagated from other units. We will discuss more about this with concrete examples in § 3.2.
@@
    Static Analyzer (§ 3.3): We collect key dependency data (e.g., call graph, dataflow analysis) using traditional program analysis techniques. Since our translation process is incremental, the program changes at every translation step, and the static analyzer must be re-run each time.
@@
    Translation (§ 3.4): The translation prompt is composed of the code component to be translated, augmented with information derived from static analysis. This information describes the surrounding context of the code segment and its dependencies, which guides the LLM in translating the unsafe Rust code to safe Rust.
@@
    Validator (§ 3.4): This component acts as a regression testing framework, validating the translation of each unit by first compiling the new code, and then running system-level tests provided along with the original C code. The safe Rust components replace their unsafe counterparts progressively, allowing both safe and unsafe Rust to coexist during validation.
@@
    Repair (§ 3.4): After a unit is translated, if there an error in validation, then the repair component of our system collects feedback from compilation and execution. Then, it prompts the LLM with this feedback and asks it to re-generate the translation of the unit.
For each translation unit, the cycle of Validation and Repair continues until the validation succeeds, or we exceed $N$ attempts (where $N$ is a parameter of our system). If we are unable to translate a unit within $N$ attempts, we leave it as its original unsafe Rust version, thus maintaining overall correctness.
In this manner, we go through the entire unsafe Rust program, processing each translation unit in order. Once we translate (or attempt to translate) a unit, we do not come back to it later. In the remainder of this section, we describe each of the components of our system in detail.
@@ 3.2.1 Decomposing the program into translation units
Our goal is to translate the entire program into safe Rust; however, we are limited by the size of the LLM context window. Further, even if the input and output fit within the input context and output context, LLMs are more prone to hallucination and translation errors when the input is long. Our solution is to decompose the program into translation units of less than $L$ lines each, and translate these units individually.
@@
Figure 2: An example of the output of our decomposition algorithm as described in § 3.2.1, applied to the cat function from the cat program. Translation Unit 1 is the “root” unit that represents the function as a whole, and all other units are nested within it. The units are annotated with comments showing live-in and live-out information, which is described in § 3.3.
@@
  • Inner translation units: The remaining $n-1$ translation units are nested within each other or within the root units.
An example of this is shown in Figure 2. Note that this decomposition does not include code that is outside of functions, like structure and enum definitions. Translating these into safe Rust is a non-trivial problem, and is out of the scope of our paper.
@@ 3.2.2 Arranging translation units in an optimal order
This gives us an ordering of functions within each WCC of the call graph. The order of functions between two WCCs doesn’t matter, so we process the WCCs one at a time in an arbitrary order to create the full ordering of all functions in the program.
This gives us an ordering of functions; however each function could contain multiple translation units. As we will discuss in § 3.4, the order of translation units within a function does not matter for our current translation strategy. However, we ensure that all the translation units corresponding to a single function are contiguous in the translation order.
@@ 3.3 Static Analyzer
We need to annotate each translation unit with enough context information to enable the LLM to translate the code well. Further, each unit cannot be translated in isolation, because it could have ripple effects on other regions of the code. Our solution is to translate all these affected regions simultaneously.
@@
Figure 4 shows slices for a root translation unit and an inner translation unit, composed into an LLM prompt.
Figure 4: An example of translating code units with LLM. For the root translation unit on the left, we generate the translation of the function body plus all of its call sites simultaneously. For the inner translation unit on the right, we translate just the code contained within it. In both cases, we also allow the model to specify any additional imports it needs.
@@ 3.4 Translation, Validation and Repair
We now attempt to translate each translation unit from unsafe Rust to safe, idiomatic Rust, using an LLM. We pass each unit to an LLM, along with its context slice as described in § 3.3. We instruct the LLM to convert the translation unit to idiomatic Rust. If the unit is a root translation unit of a function and its context contains call sites, we instruct the LLM to translate the call sites too. We specify a particular format for the LLM response that allows us to easily parse it and separate it into function and call-site translations. If the unit is an inner translation unit, we ask the LLM to ensure that the live-in and live-out variables at the beginning and end of the unit are not affected. This means that translating such a translation unit has no side-effects on any other part of the program. This substantiates our claim in § 3.2.2 that the order in which we translate the units within a function does not matter.
Sometimes, the translated unit needs additional use statements to work; we allow the LLM to specify this too in its response. An example of a prompt and LLM response is shown in Figure 4.
@@ 4.1 Datasets
We contribute a benchmark dataset of 7 C programs collected from GNU coreutils [^2] https://github.com/coreutils/coreutils. These are the basic file, shell and text manipulation utilities of the GNU operating system, and include common shell commands like ls and cat. We chose coreutils because it is a large, mature C code project, with many constituent programs of varying sizes. Further, there is an ongoing community effort to manually rewrite coreutils in Rust [^3] https://github.com/uutils/coreutils, which means that a) translating coreutils to Rust is an important real-world problem, and b) future researchers using our benchmark can compare the quality of their automatically-generated idiomatic Rust code with human-written Rust code. Finally, each benchmark in coreutils comes with end-to-end system-level test cases, which have been carefully constructed using fuzzing to ensure high coverage[4].
@@
The statistics of these 7 programs are shown in Table I. In addition to these 7 programs from coreutils, we also use 10 benchmark programs from a prior work, Laertes [8]. These do not come with end-to-end test cases. Although some of the utilities have unit tests, this is not useful because we cannot run C unit tests on Rust programs.
Program # LoC # Func Max Fn Len Mean Fn Len
Coreutils split 13848 207 1634 52.95
pwd 5859 127 902 33.02
cat 7460 166 1006 33.84
truncate 7181 124 902 44.65
uniq 8299 167 902 36.26
tail 14423 266 902 43.81
head 8047 153 902 40.77
Laertes tulipindicators 44486 234 306 32.10
@@
TABLE I: The statistics of the 7 programs in our coreutils benchmark dataset, plus the 10 programs from Laertes.

### v2

@@ 1 Introduction
Another challenge when applying LLMs for code translation is the length of input and/or output code. LLMs operate on a fixed context window, which means that they can only process inputs and generate outputs that are less than a certain predefined length. A real-world code project with dozens of files and hundreds of lines per file may not fit within this window. Further, even if the entire project does fit, it is known that LLMs tend to produce more errors as the length of the input or output increases [18]. For these reasons, most LLM-based code translation research has focused on programs that are relatively short standalone files, such as those used in competitive programming contests.
@@
Summary of Results: We evaluate on two datasets - one contributed by us, and one from a prior work, Laertes [7].
@@
    The ecosystem of static analysis tools for Rust is rather limited, and so we had to implement most of our analyses as Rust compiler passes, directly using the Rust compiler’s intermediate representations. This lays the groundwork to build future Rust static analysis tools. For instance, we build program slicing and callgraph analysis algorithms, which are quintessential static analysis techniques. We make all of this code publicly available on GitHub [^1] https://github.com/vikramnitin9/c2saferrust.
@@ 2.1 Safety and Unsafety in Rust
Rust is a systems programming language designed to offer memory safety guarantees while providing performance comparable to C and C++. In contrast, C and C++ are widely used but prone to memory errors and security vulnerabilities [6, 13, 23]. Rust achieves safety through a strict ownership model: each value in memory has a single owner, and deallocation occurs automatically when the owner goes out of scope. Values may be borrowed by other variables, but such borrows are annotated with lifetimes, which the compiler tracks to ensure validity. A common pattern is for functions to borrow values from callers for the duration of the function and return ownership afterward.
Rust’s type system distinguishes between immutable (&T) and mutable (&mut T) borrows, enforcing exclusivity rules to prevent data races. Multiple immutable borrows may coexist, as they do not modify state, but a mutable borrow requires exclusive access, disallowing any simultaneous borrows. The borrow checker ensures statically that these constraints hold, resulting in memory safety without runtime overhead [27].
While these guarantees eliminate entire classes of bugs, they also make certain programming tasks more difficult. In particular, designing data structures that rely on shared ownership or cyclic references (such as graphs) becomes non-trivial, and careful design is needed to avoid lifetime conflicts. Safety enforcement may also introduce bounds checks or borrow checking overhead and complicate interactions with low-level system components or foreign libraries [3]. To support such cases, Rust exposes the unsafe keyword, permitting raw pointer manipulation, mutable static memory, and calls to unsafe functions (including FFI and compiler intrinsics) [26]. Unsafe enables performance-critical or low-level operations but transfers responsibility for correctness from the compiler to the programmer.
The option to selectively use unsafe while retaining strong guarantees elsewhere has driven efforts to port legacy C and C++ systems into Rust, aiming to reduce vulnerability surfaces associated with manual memory management [23]. However, this process is often labor-intensive. The permissive memory model in C—with raw pointers, unchecked array access, and no concept of ownership—makes straightforward translation into Rust’s stricter semantics challenging and error-prone [14].
@@ 2.2 Translating C to Rust using Transpilation
A common approach to porting C code is the use of source-to-source compilers, or transpilers, which aim to transform code from one language into another while preserving semantics. Tools such as C2Rust [9] translate C constructs into Rust systematically and can automatically produce compilable Rust code.
Challenges.  Despite providing a starting point, transpilation often produces code that compiles but does not fully benefit from Rust’s safety guarantees. Transpilers map raw pointers to *const and *mut, frequently requiring unsafe dereferencing, and global variables in C typically become static mut, which also demands unsafe access. Common C idioms such as pointer arithmetic, manual memory management, or type reinterpretation via transmute reintroduce the possibility of undefined behavior. Moreover, transpiled code often retains low-level malloc/free semantics rather than adopting idiomatic Rust abstractions like Box, Rc, or Arc, increasing reliance on unsafe and complicating verification of memory safety.
@@ 2.3 Programming language translation with Large Language Models (LLMs)
Large Language Models offer an emerging alternative for code translation. Their ability to generate code with contextual reasoning allows them to restructure programs more flexibly than rule-based transpilers, and they may incorporate safe coding idioms learned from training corpora [17, 29, 12]. LLMs have shown promising results on structured benchmarks such as CodeNet [20], Avatar [1], and EvalPlus [15], where problems are self-contained and relatively small.
However, studies demonstrate that translating real-world codebases remains difficult. LLMs often struggle to maintain semantic equivalence across multi-file projects, preserve API interactions, or reason about complex type and memory relationships [12]. These limitations manifest as type mismatches, missing edge cases, broken dependencies, or unsafe substitutions. As a result, improving robustness and safety in LLM-based translation remains an active research direction [17].
@@ 3.1 Overview
C2SaferRust is a neuro-symbolic framework for translating C programs to their safer Rust counterparts, focusing on preserving the intent of the original code. C2SaferRust is designed to operate on complex, real-world C projects, often comprising thousands of lines of code and tens of interdependent files. The translation process prioritizes functional correctness, ensuring that the resulting Rust code behaves as expected under the test cases provided by the user.
We leverage on-demand static analysis and LLMs for translation, and differential testing for validating the translated code. However, given the size of typical C project, directly translating an entire C project to Rust in one step is infeasible. We need to come-up with a strategy to translate and validate smaller pieces of code in a step-by-step manner. Thus, we need a framework that allows partial C and partial (translated) Rust code to coexist, as we need to validate the accuracy of the translation by running this hybrid code. To this end, we adopted the C2Rust framework, which transpiles a given C code repository into an equivalent unsafe Rust repository. While the C2Rust-transpiled code retains most of C’s ‘unsafe‘ properties and original intent, it effectively acts as a Rust wrapper, enabling the unsafe code to execute within a Rust environment. Our translation process begins from this starting point, progressively transforming the unsafe Rust into safer Rust.
This unsafe Rust code is then divided into smaller units, called translation units, using program graphs extracted through symbolic reasoning. Each unit is translated into safer Rust using an LLM, which generates safer code while symbolic reasoning ensures accuracy. Compilation and test cases identify errors, prompting targeted analysis to refine context and guide effective repairs. This iterative approach balances efficiency and accuracy, progressively enhancing translation quality to suit the structure and requirements of each project. An overview of the system is shown in Figure˜1. C2SaferRust consists of the following components:
@@
    Translation Orchestrator (§˜3.2): This component, aided by static analysis, orchestrates translation. It breaks the project into smaller translation units that fit within the LLM’s context. It then selects translation units in the optimal order for translation. Translation units are not independent - when one unit is translated, the side effects have to be propagated to all other affected units. The optimal order for translation is moving up the chain of propagated side effects, i.e., translating a unit before translating the units to which its side effects are propagated. This ensures that while the LLM is translating a unit, it can also assimilate the side effects propagated from other units. We will discuss more about this with concrete examples in §˜3.2.
@@
    Static Analyzer (§˜3.3): We collect key dependency data (e.g., call graph, dataflow analysis) using traditional program analysis techniques. Since our translation process is incremental, the program changes at every translation step, and the static analyzer must be re-run each time.
@@
    Translation (§˜3.4): The translation prompt is composed of the code component to be translated, augmented with information derived from static analysis. This information describes the surrounding context of the code segment and its dependencies, which guides the LLM in translating the unsafe Rust code to safe Rust.
@@
    Validator (§˜3.4): This component acts as a regression testing framework, validating the translation of each unit by first compiling the new code, and then running system-level tests provided with the original C code. The safe Rust components replace their unsafe counterparts progressively, allowing safe and unsafe Rust to coexist during validation.
@@
    Repair (§˜3.4): After a unit is translated, if there an error in validation, then the repair component of our system collects feedback from compilation and execution. Then, it prompts the LLM with this feedback and asks it to re-generate the translation of the unit.
For each translation unit, the cycle of Validation and Repair continues until the validation succeeds, or we exceed $N$ attempts (where $N$ is a parameter of our system). If we are unable to translate a unit within $N$ attempts, we leave it as its original unsafe Rust version, thus maintaining overall correctness. In this manner, we go through the entire unsafe Rust program, processing each translation unit in order. Once we translate (or attempt to translate) a unit, we do not come back to it later. In the remainder of this section, we describe each of the components of our system in detail.
@@ 3.2.1 Decomposing the program into translation units
Our goal is to translate the entire program into safe Rust; however, we are limited by the size of the LLM context window. Further, even if the input and output fit within the context window, LLMs are more prone to hallucinations and translation errors when the input is long. Our solution is to decompose the program into translation units of less than $L$ lines each, and translate these units individually.
@@
Figure 2: An example of the output of our decomposition algorithm as described in §˜3.2.1, applied to the cat function from the cat program. Translation Unit 1 is the “root” unit that represents the function as a whole, and all other units are nested within it. The units are annotated with comments showing live-in and live-out information, which is described in §˜3.3.
@@
  • Inner translation units: The remaining n1n-1 translation units are nested within each other or within the root units.
An example of this is shown in Figure˜2. Note that this decomposition does not include code that is outside of functions, like structure and enum definitions. Translating these into safe Rust is a non-trivial problem, and is out of the scope of our paper.
@@ 3.2.2 Arranging translation units in an optimal order
This gives us an ordering of functions within each WCC of the call graph. The order of functions between two WCCs doesn’t matter, so we process the WCCs one at a time in an arbitrary order to create the full ordering of all functions in the program. Further, each function could contain multiple translation units. As we will discuss in §˜3.4, the order of translation units within a function does not matter for our current translation strategy. However, we ensure that all the translation units corresponding to a single function are contiguous in the translation order.
@@ 3.3 Static Analyzer
We need to annotate each translation unit with enough context to enable the LLM to translate the code well. Further, each unit cannot be translated in isolation, because it could have ripple effects on other regions of the code. Our solution is to translate all these affected regions simultaneously.
@@
Figure˜4 shows slices for a root translation unit and an inner translation unit, composed into an LLM prompt.
Figure 4: Translating units with an LLM. For the root translation unit (left), we generate the translation of the function body plus all of its call sites simultaneously. For the inner translation unit (right), we translate just the code contained within it. In both cases, we also allow the model to specify any imports it needs. Full prompts are in the Appendix.
@@ 3.4 Translation, Validation and Repair
We now attempt to translate each translation unit from unsafe Rust to safe, idiomatic Rust, using an LLM. We pass each unit to an LLM, along with its context slice as described in §˜3.3. We instruct the LLM to convert the translation unit to idiomatic Rust. If the unit is a root translation unit of a function and its context contains call sites, we instruct the LLM to translate the call sites too. We specify a particular format for the LLM response that allows us to easily parse it and separate it into function and call-site translations. If the unit is an inner translation unit, we ask the LLM to ensure that the live-in and live-out variables at the beginning and end of the unit are not affected. This means that translating such a translation unit has no side-effects on any other part of the program. This substantiates our claim in §˜3.2.2 that the order in which we translate the units within a function does not matter.
Sometimes, the translated unit needs additional use statements to work; we allow the LLM to specify this too in its response. An example of a prompt and LLM response is shown in Figure˜4.
@@ 4.1 Datasets
We contribute a benchmark dataset of 7 C programs collected from GNU coreutils [^2] https://github.com/coreutils/coreutils. These are the basic file, shell and text manipulation utilities of the GNU OS, and include common shell commands like ls and cat. We chose coreutils because it is a large, mature C code project, with many constituent programs of varying sizes. Further, there is an ongoing community effort to manually rewrite coreutils in Rust [^3] https://github.com/uutils/coreutils, which means that a) translating coreutils to Rust is an important real-world problem, and b) future researchers using our benchmark can compare the quality of their automatically-generated idiomatic Rust code with human-written Rust code. Finally, each benchmark in coreutils comes with end-to-end system-level test cases, which have been constructed using fuzzing to ensure high coverage [^4] https://www.pixelbeat.org/docs/coreutils-testing.html.
@@
The statistics of these 7 programs are shown in Table˜I. In addition to these 7 programs from coreutils, we also use benchmarks from a prior work, Laertes [7]. Since evaluating on each additional benchmark incurs a time cost, we selected a subset of 10 out of their 17 benchmarks, which we believe is sufficient for the purposes of a fair evaluation. These benchmarks do not come with end-to-end test cases. Although some of the utilities have unit tests, this is not useful because we cannot run C unit tests on Rust programs.
Program # LoC # Fn
# Cov.
Fn
Max
Fn Len
Mean
Fn Len
Coreutils
split 13848 207 73 1634 52.95
pwd 5859 127 16 902 33.02
cat 7460 166 37 1006 33.84
truncate 7181 124 33 902 44.65
uniq 8299 167 34 902 36.26
tail 14423 266 76 902 43.81
head 8047 153 45 902 40.77
Laertes
tulipindicators 44486 234 306 32.10
@@
TABLE I: The statistics of the 7 programs in our coreutils benchmark dataset, plus the 10 programs from Laertes. “# Cov. Fn” is the number of functions covered by tests.